OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sec2x - 2 = tan2x

OpenStudy (anonymous):

do you mean Find all solutions sec^2x + tan^2x = 3 in interval [0,2pi)?

OpenStudy (anonymous):

yes. Sorry about that.

OpenStudy (anonymous):

sec²(x) + tan²(x) = 3 Pythagorean trig identity: sec²(x) + (sec²(x) - 1) = 3 Combine like terms on left: 2sec²(x) - 1 = 3 Add 1 to both sides: 2sec²(x) = 4 Divide both sides by 2: sec²(x) = 2 Take the square root of both sides: sec(x) = ±√2 Put in terms of cosine: 1/cos(x) = ±√2 Multiply both sides by ±cos(x)/√2: cos(x) = ±1/√2 Multiply right-hand side by √2/√2: cos(x) = ±√2/2 Use the unit circle to evaluate: x = π/4, 3π/4, 5π/4, 7π/4